# Grasping the Essentials: Tailoring Large Language Models for Zero-Shot Relation Extraction

### v1

@@ 4.1 Evaluation Dataset Construction
Dataset #test rels #test #unlabel
DefOn-FewRel 14 9,798 100,000
DefOn-ReTACRED 11 2,148 100,000
Table 1: Brief statistics about datasets constructed.
Model DefOn-FewRel DefOn-ReTACRED
[…]
@@
Random Guess 7.14 49.76 12.49 9.10 50.02 15.37
GPT-3.5 50.79 69.76 54.23 49.58 38.48 42.52
RoBERTa NLI 39.67 94.34 49.40 26.94 97.55 40.64
ZS-BERT* 41.00 40.51 40.73 20.20 17.70 18.81
RelationPrompt* 75.18 66.08 70.34 51.67 51.26 51.40
RelationPrompt (NoGen)* 0.00 0.00 0.00 2.98 0.83 1.25
RE-Matching* 76.47 71.80 74.05 54.43 50.22 52.16
REPaL (Ours) 78.19 82.25 77.93 68.47 80.52 68.42
Table 2: Evaluation results of relation extraction models under zero-shot and few-shot settings. Zero-shot models marked by * are trained on seen relation instances and require all negative test relations to be known. REPaL is trained on 15p15n initial instances and further trained with additional 15p15n follow-up instances.
We first construct datasets for model evaluation on the Definition Only Zero-Shot Relation Extraction task. Our constructed datasets are based on FewRel (Han et al., 2018b) and ReTACRED (Zhang et al., 2017; Stoica et al., 2021) respectively. As FewRel’s relations are from WikiData’s  [^3] WikiData main page properties, we query the property definitions and slightly modify them to become complete sentences with entity placeholders [ENT0] and [ENT1] (corresponding to $E_{0}$ and $E_{1}$ as in definition notation $d(E_{0},E_{1})$). As TACRED/ReTACRED’s data are based on 2009-2014 TAC KBP (Zhang et al., 2017), we annotate relation definitions based on the guidelines (especially the slot definitions  [^4] TAC KBP guidelines) as well as related adjustments introduced in ReTACRED paper (Stoica et al., 2021).
The evaluation relations are chosen based on the following general principles: (1) The selected relation should has abundant instances for evaluation (generally, we set a minimum requirement of 200 instances); (2) The annotated instances should follow the officially documented relation definitions and the sentence should be sufficient to deduce the target relations without external knowledge; and (3) The selected relations should be semantically disjoint without ambiguity. These principles lead to the brief dataset statistics in Table 1.
To construct unlabeled corpus, we down-sample the FewRel unlabeled corpora (from WikiData) (Han et al., 2018b; Gao et al., 2020) to 10,000 instances. For ReTACRED (Zhang et al., 2017; Stoica et al., 2021), our constructed unlabeled corpora includes all unselected instances and sampled instances from FewRel unlabeled corpus so that the unlabeled corpus size also reaches 10,000. More details on the two constructed datasets can be found in Appendix C.
@@ 4.2 Baseline Methods
We compare our method with the following baselines under Fully-Supervised, Few-Shot, and Zero-Shot settings: (1) Random Guess: gives uniform random binary predictions. (2) GPT-3.5 (Brown et al., 2020; Ouyang et al., 2022): uses gpt-3.5-turbo-0125 model and formulates RE as a double choice problem, answering whether two entity mentions in each test instance follow the given target relation definition. GPT-3.5 ICL on DefOn-ReTACRED uses randomly sampled 10 positive and 10 negative (abbreviated as 10p10n) gold instances as demonstrations while adopts 5p5n for DefOn-FewRel (better than 10p10n and 2p2n). (3) RoBERTa NLI (Devlin et al., 2018; Liu et al., 2019): our adopted SLM RE backbone model (Sec. 3.2). It uses 100p100n for DefOn-FewRel and 70p70n for DefOn-ReTACRED for Fully-Supervised setting. (4) ZS-BERT (Chen and Li, 2021): a Siamese-based zero-shot RE framework trained on seen labeled instances and gives prediction by nearest neighbour search comparing encoded sentence representations and relation definition representations. (5) RelationPrompt (Chia et al., 2022): a Seq2Seq-based zero-shot RE framework that trains GPT-2 (Radford et al., 2019) to generate relation instances conditioned on relation names and train BART (Lewis et al., 2020) to generate the extracted relation triple on seen data. For unseen data, it fine-tune BART on synthetic data generated by trained GPT and then predicts. (6) RE-Matching (Zhao et al., 2023): a Siamese-based zero-shot RE model that encodes entity types and relation definitions for each relation on seen data and conducts nearest neighbour search for prediction on unseen data. We also test on RE-Matching (NoGen) that removes the data generator.
Note that ZS-BERT, RelationPrompt and RE-Matching require being trained on labeled seen data and their predictions on unseen data are conducted in a multi-way classification manner where information of all negative relations is leveraged. These three baselines are trained on 61 FewRel relations with another 5 FewRel relations in dev set (each relation has 700 instances).
@@ 4.3 Experiment Setup
Following our Definition Only Zero-Shot Relation Extraction setting, evaluation is conducted similar to cross validation where each iteration one test relation is treated as positive target relation and remaining test relations serve as negative test relations (our setting assumes all negative relations are unknown). Macro $F_{1}$ is computed by averaging across all iterations. Reported results in this work are the average over random seeds 1, 2, and 3. See Appendix E for implementation details.
@@ 5.1 Main Results
The main results are shown in Table 2. REPaL achieves much better performance compared with all few-shot and zero-shot baselines with large margins. As the strong baselines are pre-trained on FewRel data, not ReTACRED data, REPaL’s advance is much larger in DefOn-ReTACRED. This shows the effectiveness of our method in low-resource settings and the robustness across different benchmarks as we do not rely on any benchmark-specific supervision. Compared with fully-supervised baselines, it shows there is still room for improvement. This is likely due to our task’s assumption on unknown negative test relation space which is normally avoided in mainstream zero-shot RE works.
@@ 5.2 Ablation Study
[…]
REPaL
   RoBERTa NLI 39.67 94.34 49.40
   w/ P_init 67.22 90.11 73.76
  +P_feedback 72.56 87.79 76.84
   +N_feedback 78.19 82.25 77.93
Table 3: Ablation results with 15p15n as initial setting. P_init, P_feedback, and N_feedback denote initial positive generation, feedback-driven follow-up positive generation and negative generation respectively.
To investigate the effectiveness of our framework design, we conduct ablation studies with results shown in Table 3. The initial seed generation brings considerable performance advance which shows LLM’s power of domain adaptation is fully leveraged given a clear relation definition. Based on experiments on Sec. 5.4, the best $F_{1}$ achieved by a single seed among all initial seed generation setups is 75.92 (p=60, n=60, random seed=3). So by adopting our feedback-driven follow-up generation design, in total p=30 & n=30 surpasses this upper bound with approximately 2 points.
Furthermore, the negative follow-up instance generation further boosts the precision score, demonstrating its effectiveness on rectifying SLM’s bias in distinguishing positive and negative relations.

### v2

@@ 4.1 Evaluation Dataset Construction
Model DefOn-FewRel DefOn-Wiki-ZSL
Precision Recall F1 Macro-F1 Precision Recall F1 Macro-F1
@@
   Random Guess 7.14 50.77 12.52 - 6.67 51.01 11.67 -
   GPT-3.5 55.09 61.81 53.94 - 42.64 45.70 39.60 -
   RE as QA 84.57 73.42 74.34 84.67 75.72 61.48 60.89 70.27
   RoBERTa NLI 48.79 81.17 55.07 - 36.23 62.65 35.28 -
   ZS-BERT 35.91 35.70 - 35.78 37.11 33.85 - 35.26
   RelationPrompt 74.39 66.89 - 67.78 71.89 60.50 - 61.08
   RE-Matching 77.49 72.95 - 75.11 73.10 68.99 - 70.97
   REPaL (w GPT-4o mini) 73.27 73.75 70.46 90.01 65.67 43.53 44.78 78.47
   REPaL (w GPT-4o) 78.86 77.28 74.61 91.71 68.98 47.63 47.80 80.96
Table 1: Evaluation results of relation extraction models under zero-shot and fully-supervised settings. REPaL is trained on 15p15n initial instances to acquire SLM inference feedback and then trained with additional 15p15n follow-up instances (30p30n in total). We show the results with different backbone synthesis LLMs (GPT-4o and GPT-4o mini). Note that the results for GPT-3.5 and RE as QA baselines on the Definition Only Zero-Shot Relation Extraction setting are based on the down-sampled subsets with 30 samples/relation to reduce inference cost.
We first construct datasets for model evaluation on the Definition Only Zero-Shot Relation Extraction task. Our constructed datasets are based on FewRel (Han et al., 2018b) and Wiki-ZSL (Chen and Li, 2021) respectively [^4] An ideal test set for our setting should follow: (1) annotations should follow the officially documented relation definitions and the sentences should be sufficient to deduce the target relations without external knowledge; (2) relations are better semantically disjoint without ambiguity so that we do not need to manually adjust binary test labels for overlapped relations in each test iteration. These principles also apply to existing multi-way classification RE datasets. However, we leave these for future dataset related work.. The public portion of FewRel contains 80 relations, each comes with 700 instances annotated by crowd workers. Wiki-ZSL consists of 113 relations with total 93,383 instances obtained by distant supervision. As their relations are from WikiData’s [^5] WikiData main page properties, we query the property definitions and slightly modify them to become complete sentences with entity placeholders [ENT0] and [ENT1] (corresponding to $E_{0}$ and $E_{1}$ as in definition notation $d(E_{0},E_{1})$). For evaluation, we sample 5 groups of 14 FewRel relations and 3 groups of 15 Wiki-ZSL relations as test sets.
To construct unlabeled corpus, for each group of the test set, we down-sample 10,000 instances from the unlabeled corpora used by Gao et al. for few-shot relation learning which contains 744 distantly supervised relations (labels are not used in this work) and in total 899,996 instances (2020). The final test sets with the corpus are denoted as DefOn-FewRel and Defon-Wiki-ZSL respectively.
@@ 4.2 Baseline Methods
We compare our method with the following baselines under Fully-Supervised and Zero-Shot settings: (1) Random Guess: gives uniformly random binary predictions. (2) GPT-3.5 (Brown et al., 2020; Ouyang et al., 2022): uses gpt-3.5-turbo-0125 model and formulates RE as a double choice problem, answering whether two entity mentions in each test instance follow the given target relation definition. (3) RE as QA: similar to QA4RE (Zhang et al., 2023), we design more QA-like prompt templates based on the relation definitions and gpt-3.5-turbo-0125 model to formulate RE as a multiple choice problem (double choice for Definition Only Zero-Shot Relation Extraction task and multi-choice for traditional zero-shot RE). (4) RoBERTa NLI (Devlin et al., 2018; Liu et al., 2019): our adopted SLM RE backbone model (Sec. 3.2). It adopts 100 gold positive and 100 gold negative instances for each relation under the Fully-Supervised setting. (5) ZS-BERT (Chen and Li, 2021): a Siamese-based zero-shot RE framework trained on seen labeled instances and gives prediction by nearest neighbour search comparing encoded sentence representations and relation definition representations. (6) RelationPrompt (Chia et al., 2022): a Seq2Seq-based zero-shot RE framework that trains GPT-2 (Radford et al., 2019) to generate relation instances conditioned on relation names and train BART (Lewis et al., 2020) to generate the extracted relation triple on seen data. For unseen data, it fine-tune BART on synthetic data generated by trained GPT and then predicts. (7) RE-Matching (Zhao et al., 2023): a Siamese-based zero-shot RE model that encodes entity types and relation definitions for each relation on seen data and conducts nearest neighbour search for prediction on unseen data.
Note that ZS-BERT, RelationPrompt and RE-Matching require being trained on labeled seen data and their predictions on unseen data need to be conducted in a multi-way classification manner where information of all relations is leveraged. These three baselines are trained on the relation instances not in the test set while leaving 5 relations’ instances for dev set.
@@ 4.3 Experiment Setup
Following our Definition Only Zero-Shot Relation Extraction setting, evaluation is conducted similar to cross validation where each iteration one test relation is treated as the gold positive target relation and remaining test relations serve as gold negative test relations (our setting assumes all negative relations are unknown in terms of relation definitions and any distribution information). Each test iteration is a binary classification problem with precision, recall and F1 calculated. Table 1 shows the main results where Precision, Recall and F1 indicates the corresponding binary classification metric scores averaged across all test iterations. If the cell for F1 column is empty, then the Precision and Recall represents the Macro Precision and the Macro Recall corresponding to Macro F1. Since ZS-BERT, RelationPrompt and RE-Matching require the traditional zero-shot RE setting where multi-way classification is assumed, we further train the KnowPrompt (Chen et al., 2022b) multi-way RE classifier with all the positive instances of test relations synthesized by REPaL in the Definition Only Zero-Shot Relation Extraction setting for comparison. This leads to the Macro-F1 column. See Appendix F for further implementation details and see Appendix A for details of our definition only zero-shot RE setting.
@@ 5.1 Main Results
The main results are shown in Table 1. REPaL achieves generally better performance compared with zero-shot baselines with large margins for both evaluation datasets. This shows the effectiveness of our method in low-resource settings and the robustness across different benchmarks.
The advance of REPaL is slightly lower in DefOn-Wiki-ZSL compared to DefOn-FewRel which is because Wiki-ZSL is much more noisy due to distant supervised annotations. By comparing the absolute values of F1 scores derived for our zero-shot RE setting with the Macro-F1 scores derived for traditional zero-shot RE setting, it can be concluded that our definition only zero-shot setting is much more challenging. Once we break the assumption of unknown negative relation space, models can take shortcuts to distinguish different relations without actually comprehending the relations.
Compared with fully-supervised baselines, it shows there is still room for improvement. This is related to our task’s assumption on unknown negative test relation space which is normally overlooked in mainstream zero-shot RE works. Note that, on our Definition Only Zero-Shot RE setting, RE as QA achieves better performance on Defon-Wiki-ZSL but lower on traditional zero-shot RE setting. This shows that LLMs are capable of judging whether an instance follows a certain relation if given a clear relation definition. However, they are less competent for dealing with multiple relations. Meanwhile, inference with LLMs are much more costly as the RE as QA (w/ GPT-3.5) would cost around \$260 for one DefOn-FewRel split on our setting if without down-sampling. For traditional setting, the RE as QA would cost around \$2.3. REPaL, in contrast, even with GPT-4-0125-preview, costs around \$3.7 for generating 30p30n train and 30p30n dev examples [^6] Approximated by multiplying the number of input and output tokens of GPT-4o with the price of GPT-4-0125-preview.. The generating instances are further reused for both settings.
@@ 5.2 Ablation Study
Precision Recall F1
REPaL 78.86 77.28 74.61
   - N_feedback 76.74 77.47 73.04
    - P_feedback 73.59 75.90 70.03
     - P_init 48.79 81.17 55.07
Table 2: Ablation results of REPaL. P_init, P_feedback, and N_feedback denote initial positive generation, feedback-driven follow-up positive generation and negative generation respectively. The ablations are based on REPaL w GPT-4o.
To investigate the effectiveness of our framework design, we conduct ablation studies with results shown in Table 2. The initial seed generation brings considerable performance advance which shows LLM’s power of domain adaptation is fully leveraged given a clear relation definition. Compared to the results shown in Sec. 5.3 where we generate 30p30n instances all at once, REPaL achieves better recall and F1 both with and without follow-up negative instance generation. This indicates the importance of feedback-driven generation design.
Furthermore, the negative follow-up instance generation further boosts the precision, demonstrating its effectiveness on rectifying SLM’s bias in distinguishing positive and negative relations.